Test
PRE – TEST
Preguntas del test (3)
1The probability that a sum of seven appears in a single toss of fair dice is
- A\[\frac{1}{6}\]
- B\[\frac{5}{6}\]
- C\[\frac{1}{36}\]
- D0
- E\[\frac{1}{10}\]
2The probability of drawing three aces from a pack of cards is ______________________ the probability of drawing the 8 of hearts, then the three of diamond, then the three of diamonds, then the queen of cluds.
- Agreater than
- Bless than
- Cequal to
3Events M and Q are mutually exclusive. Which of the following is TRUE?
- AM and Q are also independent
- BP(M)P(Q) = P(M and Q)
- CP(M or Q) = P(M) + P(Q)
- DP(M or Q) = 1 – P(M and Q)
Información
- Descripción
- Test sobre PRE – TEST, creado por gdumitra el 29/01/2015.
- Autor
- gdumitra
- Idioma
- English
- Visitas
- 5
- Valoración
- Aún sin valoraciones Valora este test